Solve each equation
5-10|8n+5|=-45

Answer by stanbon(75887)

5-10|8n+5|=-45
-----
-10|8n+5| = -50
----
|8n+5| = 5
----
8n + 5 = 5 or 8n+5 = -5
8n = 0 or 8n = -10
n = 0 or n = -5/4
======================
Cheers,
Stan H.
==============